# The Open Journal of Astrophysics

### v1

@@ 2.3 Training Specifics
We set the base channel of our U-net to 64, with channel multipliers at different resolutions set to [1, 1, 2, 2]. For the discriminator network, the base channel is 32 and the channel multipliers at each resolution are [1, 2, 4, 8]. The Eulerian field, sampled on meshes of side length 1024 cells, is concatenated with a 6D field of particle displacements and velocities, serving as the input for our discriminator. Data augmentation techniques are applied, including random flips along three different axes and random shifts of the chunk crop position. We use Rectified Adam (Liu et al., 2019) in training to avoid identified convergence problem suffered by the Adam optimizer. We set parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$ which is the default in Pytorch (Paszke et al., 2019). The batch size is set to 1 which is the largest size possible given our setup. Training was conducted using NVIDIA A100 GPUs.
@@ 3 Results
We now compare the outputs of our emulated SR simulations with full N-body HR runs. In our previous work on stochastic SR simulations (Li et al., 2021; Ni et al., 2021; Zhang et al., 2023), we compared the results using various summary statistics, including power spectra and halo mass functions. In the current work, because we are emulating the exact HR fields, we can perform comparisons on an individual basis rather than relying solely on statistical measures. Consequently, we supplement our analyses with additional approaches, such as visual comparisons of individual structures and cross-correlation of the individual Fourier modes between the emulated and HR runs. We also perform comparison at the halo level with the corresponding HR simulation.
3.1 Visual comparison
Figure 2: 3D visualization of the SR, Emulator, and HR dark matter density fields with boxsize $100h^{-1}{\rm Mpc}$ (first row), $20h^{-1}{\rm Mpc}$ (second row), $11h^{-1}{\rm Mpc}$ (third row) at $z=1$.
@@ 3.1 Visual comparison
For our initial validation, we visually analyze the differences between the HR fields obtained from $N$-body simulations and the SR fields of our SR model, as well as the output from our newly designed Emulator Neural Network. Figures 2 and 3 present three-dimensional visualizations of dark matter density plots, generated by projecting all particles contained within the box. The images are rendered using the gaepsi2 code. [^2] https://github.com/rainwoodman/gaepsi2 The first row shows the whole $(100h^{-1}{\rm Mpc})^{3}$ volume, centered on several massive halos, to show the large-scale structures, the second row and the third row shows a sub-volume of $(20h^{-1}{\rm Mpc})^{3}$ and $(11h^{-1}{\rm Mpc})^{3}$, respectively, to illustrate finer details and halo distributions.
The visual representation of the density fields reveals that both our SR model and emulator are capable of generating large-scale structures apearing morphologically similar to one another. Regarding the small scales, the SR model produces visually authentic features based on the constraint from the LR input, with the fine details appearing reasonable. Our Emulator refines these small-scale details to better match the HR by conditioning on the HRICs. The SR model successfully forms halos in regions where the LR input lacks resolution, while the Emulator further adjusts their positions to align with the HR simulation. As an example, as shown in the bottom row of Figure 3, the SR field generates a series of halos lying along a thin filament, and the emulator adjusts the field to align with the HR field, placing three massive halos at the center of the box.
@@ 3.2 Fourier analysis: cross-correlation of individual modes
To quantify agreement of the velocity field, we examine the momentum in Fourier space. The momentum field $p(x)$ is defined as momentum at position $x$. This is calculated using the same CIC scheme as the matter density field. The Eulerian momentum power spectrum is given by:
@@
We use the same definition as Eq. 10 for the cross-correlation. The results are plotted in Figure 4, with the color of each curve representing the redshift of its corresponding simulation. We select 5 snapshots from each simulation at redshift $z=1,0.8,0.5,0.2,0$.
@@
The results of the density power spectrum error are presented in the top row of Figure 4. The outputs of the Emulator model are considerably more accurate than those of the SR model. At $k=1\;h\;{\rm Mpc}^{-1}$, which is around the LR Nyquist frequency, the SR density field deviates from the HR model by exhibiting 7% stochasticity, whereas our emulator maintains less than 2% across all redshifts. The discrepancy between the SR and HR fields increases to over 99% stochasticity at $k=8.5h{\rm Mpc}^{-1}$, but the Emulator significantly reduces this to around 50%. The error at the HR Nyquist frequency is also reduced from 100% for SR to between 55-80% in the Emulator’s output, depending on redshift. A significant fraction of the structures are reproduced even even at $k=10\;h\;{\rm Mpc}^{-1}$ and beyond. The second and third rows show the error in the displacement and momentum power spectra, respectively. We observe significant improvements in accuracy with the Emulator model. The displacement power spectrum error in the SR simulation, which is around 60%, is reduced to less than 10% by our Emulator. Regarding the momentum power spectrum error, the Emulator reduces the SR simulation’s error from 40% to less than 5%.
@@
In the Emulator panel for the stochasticity of the density modes, we display another blue dashed line. This represents the results of the Emulator trained on a single redshift ($z=0$). This version shows a significant improvement compared to the multiple-redshift model, reducing the error at the HR Nyquist frequency to around 12%, and the error at $k=10\;h\;{\rm Mpc}^{-1}$ to only 5%. These results represent the best that the emulator is able to achieve, but we have not been unable so far to reproduce this accuracy when the emulator is trained for multiple redshifts. While our emulator in its current state is aiming for use cases currently carried out by approximate methods, it is interesting to note that a 5% error in Fourier space at $k=10\;h\;{\rm Mpc}^{-1}$ is comparable to many full N-body simulation codes (see e.g., Figure 14 of Heitmann et al. 2008).
@@
We can ask what limits the emulator’s accuracy, and how this can be improved: the movements of particles within a virialized halo are rapid and can vary significantly over time. This high frequency behavior diminishes the emulator’s ability to accurately match the positions and velocities of particles at the individual level. This task is challenging because the HRICs do not contain information within non-linear regimes. Additionally, due to memory constraints, the full simulation must be split into chunks, with the GPU handling a maximum of $128^{3}$ particles per chunk. A single large halo may span multiple chunks, meaning the emulator lacks access to all the relevant particle information. This limitation could be mitigated by using a GPU with more memory or use a more memory efficient model. Despite this, the Emulator demonstrates strong performance, effectively adjusting the small-scale details generated by the SR model starting from $k=1\;h\;{\rm Mpc}^{-1}$, achieving excellent accuracy up to the Nyquist frequency of the initial grid. It also maintains decent accuracy up to $k=10\;h\;{\rm Mpc}^{-1}$.
@@
Figure 4: Stochasticity, derived from the correlation coefficient of individual Fourier modes as a function of wavenumber. From left to right, we present results for the cross-correlation of HR (the true N-body simulation) with SR and the Emulator output. The color of each curve corresponds to the redshift, as indicated by the color bar. The shaded area shows the $1\sigma$ standard deviation from all test sets. The vertical dotted gray line is the Nyquist frequency of the LR simulation and the dashed gray line indicates the Nyquist frequency of HR simulation. The blue-dashed line in the top right panel shows the result for a version of the emulator trained on a single redshift ($z=0$, see text).
@@ 3.3 Halo catalog analysis
Dark matter, under gravitational clustering, forms high-density virialized structures and local maxima within them, known as halos and subhalos. Understanding these structures and substructures that potentially host galaxies is crucial for connecting theoretical models to observations. These halos and subhalos, identified using halo finder algorithms, play a key role in constructing mock observational catalogs from N-body simulations. We use the SUBFIND code (Springel et al., 2001) to determine the positions and masses of halos and subhalos, as well as to build merger trees. SUBFIND identifies substructures by detecting locally dense and gravitationally bound groups of particles. It starts with the Friends-of-Friends algorithm, followed by the estimation of each particle’s local density via adaptive kernel estimation.
We employ the SUBFIND algorithm on test sets of SR output, Emulator output and HR output of the $N$-body simulation and compare statistical properties of the halo and subhalo populations. Throughout this work. we examine halos and subhalos with at least 100 tracer particles in our test sets.
3.3.1 Host and subhalo abundance
The most straightforward comparison involves counting the number of halos as a function of mass. The halo mass function is defined as:
@@ 3.3.1 Host and subhalo abundance
Figure 5 and Figure 6 show the mass function for host halos and subhalos measured from the SR, Emulator, and HR field test sets across five redshifts from $z=1$ to $z=0$. The colors of the curves represent the redshift of each simulation, with shaded regions indicating the $1\sigma$ standard deviation. In Figure 5, we present host halo mass function in the top panel and its fractional error in the bottom panel. In Figure 6 we show the subhalo mass function in the same format.
@@
Figure 5: The host halo mass function at redshifts $z=1,0.8,0.5,0.2,0$ measured from the SR, Emulator and HR fields. The color of each curve denotes the redshift of simulation in test set. The shaded areas indicate $1\sigma$ standard deviation from all test sets. The upper panel shows halo mass functions and lower panel shows fractional errors.
@@
Figure 6: The subhalo mass function at redshifts $z=1,0.8,0.5,0.2,0$ calculated from the SR, Emulator and HR fields. The color of each curve denotes the redshift of simulation in test set. The shaded areas indicate $1\sigma$ standard deviation from all test sets. The upper panel shows halo mass functions and lower panel shows fractional errors.
@@
For host halos, we observe that the Emulator is slightly less accurate than the SR model within the lower mass region. particularly at $z=0$. Regarding substructures, Figure 6 shows that our Emulator model statistically improves over the SR model in the mass range covering $[10^{11},10^{13}]\,M_{\odot}$. The SR model exhibits an approximately 90% deficit in the subhalo mass function at the lower mass end, whereas our Emulator significantly reduces this error by at least 50%.
Figure 7 shows the mean occupation number of subhalos as a function of the host halo mass. The SR field has an overall offset from the HR value, with from $50\%-90\%$ fewer subhalos being predicted per halo than in the HR simulation. The emulator improves on this from the lower mass end up to $10^{13}\,M_{\odot}$, reducing the error to between 16% and 75%. As shown in Figure 8, the SR output fails to accurately predict the correct subhalos. However, our emulator model successfully predicts the correct number of subhalos with similar relative positions.
@@
Both the SR model and the emulator predict fewer subhalos within the most massive halos. We believe that this is partly due to the limited Eulerian field resolution within our discriminator. With the current setup of 512 mesh cells along each side of a $100h^{-1}{\rm Mpc}$ box, the resolution per Eulerian field "pixel" is approximately $0.2h^{-1}{\rm Mpc}$, limiting the discriminator’s ability to detect real substructures within massive halos. Increasing the resolution by a factor of $f$ would expand the Eulerian field by a factor of $f^{3}$, demanding more GPU memory. Despite these challenges, our emulator remains quite successful at predicting substructures. Across the halo mass range from $10^{11}-10^{15}\,M_{\odot}$, the mean number of subhalos per halos varies by three orders of magnitude, with the emulator providing relatively accurate results.
@@
We discover that some halos are divided between simulation chunks that are processed separately. This occurs during both the training and testing phases and causes the model to lack complete halo information. That this occurs is due to GPU memory constraints which requires the simulation to be broken up into relatively small chunks. To quantify the effect of this, we select halos where 99% of their particles are contained within the same chunk. The blue dashed line in the lower panel of the Emulator illustrates the result, showing 10% improvement in the error margin at approximately $10^{11}\,M_{\odot}$. This highlights the significance of the chunking issue in the SR+Emulator procedure. The problem could be mitigated by using a more memory-efficient model or by splitting the model sequentially across multiple GPUs for training and testing.
@@
Figure 7: The mean occupation number of subhalos vs host halo mass at redshift $z=1,0.8,0.5,0.2,0$ measured from the SR, Emulator and HR fields. The $y$ axis is the averaged number of subhalos in each host halo mass bin. The color of each curve corresponds to its redshift according to the color bar at the top. The shaded area shows the $1\sigma$ standard deviation from all test sets. The dashed blue line represents the mean occupation number at $z=0$, calculated by selecting only those halos where 99 percent of their particles lie within a single particle chunk (see text).
@@
Figure 8: Visual comparison of subhalos from SR, Emulator and HR output, centered on a massive halo with halo mass $1.1*10^{13}\,M_{\odot}$. We mark host halos with blue circles and subhalos with red circles. The radius of the circle is proportional to the virial radius. All halos shown here have mass $>100M_{DM}$.

### v2

@@ 2.3 Training Specifics
We set the base channel of our U-net to 64, with channel multipliers at different resolutions set to [1, 1, 2, 2] and attention at the last level. For the discriminator network, the base channel is 32 and the channel multipliers at each resolution are [1, 2, 4, 8]. The Eulerian field, sampled on meshes of side length 1024 cells, is concatenated with a 6D field of particle displacements and velocities, serving as the input for our discriminator. Data augmentation techniques are applied, including random flips along three different axes and random shifts of the chunk crop position. The input into the Emulator consists of overlapping chunks of size 128, with the overlap region determined by random shifts in the chunk crop position. Periodic boundary condition is applied during the random cropping. We use Rectified Adam (Liu et al., 2019) during training to avoid identified convergence problem suffered by the Adam optimizer, which may converge to bad/suspicious local optima. We set parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$ which is the default in Pytorch (Paszke et al., 2019). The batch size is set to 1 which is the largest size possible given our setup limited by the GPU memory. Training was conducted using NVIDIA A100 GPUs.
@@ 3 Results
We now compare the outputs of our Emulator model with full N-body HR runs. In our previous work on stochastic SR generations (Li et al., 2021; Ni et al., 2021; Zhang et al., 2023), we evaluated results using various summary statistics, including power spectra and halo mass functions. In this study, because we emulate exact HR fields, we are able to perform particle-level comparisons rather than relying solely on statistical measures. To complement this, we incorporate additional analyses, including visual comparisons of individual structures, evaluation of prediction errors for particle displacement and velocity, and cross-correlation of individual Fourier modes between the emulated and HR runs. Furthermore, we compare the halo-level properties of the emulated and HR simulations by analyzing merger trees and mass histories.
3.1 Direct comparison
Figure 2: 3D visualization of the SR, Emulator, and HR dark matter density fields with boxsize $100h^{-1}{\rm Mpc}$ (first row), $20h^{-1}{\rm Mpc}$ (second row), $11h^{-1}{\rm Mpc}$ (third row) at $z=1$ from same test set.
@@ 3.1 Direct comparison
For our initial validation, we perform a visual analysis of the differences between the HR fields from $N$-body simulations the SR fields generated by our SR model, and the outputs of our newly designed Emulator model. Figures 2 and 3 present three-dimensional visualizations of dark matter density plots, created by projecting all particles contained within the box. The images are rendered using the gaepsi2 code. [^2] https://github.com/rainwoodman/gaepsi2 The first row shows the whole $(100h^{-1}{\rm Mpc})^{3}$ volume, centered on several massive halos, highlighting the large-scale structures. The second and the third rows focus a sub-volume of $(20h^{-1}{\rm Mpc})^{3}$ and $(11h^{-1}{\rm Mpc})^{3}$, respectively, to illustrate finer details and halo distributions.
The visual representation of the density fields reveals that both our SR and Emulator models are capable of generating large-scale structures appearing morphologically similar to one another. On smaller scales, the SR model produces visually authentic features constrained by the LR input, with the fine details appearing reasonable. Our Emulator further refines these small-scale details to better match the HR by conditioning on the HRICs. While the SR model successfully forms halos in regions where the LR input lacks resolution, the Emulator adjusts their positions to align with the HR simulation. For instance, as shown in the bottom row of Figure 3, the SR field generates a series of halos lying along a thin filament, the emulator adjusts the field to align with the HR field, placing three massive halos at the center of the box.
Figure 4: Error distribution of displacement and velocity at $z=0$ and $z=1$. The dashed line indicates the 95th percentile error of the Emulator and the SR output.
We begin by validating our results using particle-level error. Figure 4 shows the distribution of root mean squared error distributions for displacement and velocity at $z=0$ and $z=1$. The Emulator’s error is shown in blue, while the SR error is shown in red. The Emulator model significantly reduces displacement error, with the blue dashed line indicating the 95th percentile error. At $z=0$, 95 % of particles from the Emulator have a displacement error below approximately $1.3h^{-1}{\rm Mpc}$, while at $z=1$, this threshold decreases around $0.8h^{-1}{\rm Mpc}$. In comparison, the red dashed line indicates the 95th percentile displacement error for the SR output, which is significantly higher than that of the Emulator. Specifically, the SR 95th percentile error is approximately $2.4h^{-1}{\rm Mpc}$ at $z=0$ and $1.5h^{-1}{\rm Mpc}$ at $z=1$, roughly double the Emulator’s threshold. Similarly, velocity error distributions also show substantial improvement. The Emulator’s 95th percentile velocity error is about $730\,{\rm km}\,{\rm s}^{-1}$ at $z=0$ and $460\,{\rm km}\,{\rm s}^{-1}$ at $z=1$, whereas the SR velocity error is around $1000\,{\rm km}\,{\rm s}^{-1}$ $z=0$ and $630\,{\rm km}\,{\rm s}^{-1}$ at $z=1$. These results prove that our Emulator is capable of reducing the stochasticity in the SR model’s output, impoving the accuracy significantly. The enhanced alignment of particle positions and velocities with the HR simulation confirms that the Emulator successfully utilize the information from the HRICs to guide the adjustments of the stochastic SR model’s output.
@@ 3.2 Fourier analysis: cross-correlation of individual modes
To quantify agreement of the velocity field, we examine the momentum in Fourier space. The momentum field $p(x)$ is defined as momentum at position $x$. This is calculated using the same Cloud In Cell (CIC) scheme as the matter density field. The Eulerian momentum power spectrum is given by:
@@
We use the same definition as Eq. 10 for the cross-correlation. The results are plotted in Figure 5, with the color of each curve representing the redshift of its corresponding simulation. We select 5 snapshots from simulation at redshift $z=1,0.8,0.5,0.2,0$.
@@
The results of the density power spectrum error are presented in the top row of Figure 5. The outputs of the Emulator model are considerably more accurate than those of the SR model. At $k=1\;h\;{\rm Mpc}^{-1}$, near the LR Nyquist frequency, the SR density field deviates from the HR model with a stochasticity of approximately 7%, whereas our emulator maintains less than 2% across all redshifts. The discrepancy between the SR and HR fields increases to over 99% stochasticity at $k=8.5h{\rm Mpc}^{-1}$, but the Emulator reduces this to around 50%. The error at the HR Nyquist frequency is also reduced from 100% for SR to between 55% and 80% for the Emulator, depending on redshift. Notably, a significant fraction of the structures are reproduced even at $k=10\;h\;{\rm Mpc}^{-1}$ and beyond.
The second and third rows show the error in the displacement and momentum power spectra, respectively. We observe significant improvements in accuracy with the Emulator model. The displacement power spectrum error in the SR simulation, which is around 60% near the LR Nyquist frequency, is reduced to less than 10% by our Emulator. Regarding the momentum power spectrum error, the Emulator reduces the SR simulation’s error from 40% to less than 5% near the LR Nyquist frequency.
@@
We also trained our Emulator at a single redshift. The blue dashed line and red dashed line in the Emulator panels represent the single-redshift Emulator at $z=0$ and $z=1$, respectively. The single-redshift version demonstrates comparable performance to the multi-redshift model, indicating that our multi-redshift approach experiences almost no performance loss compared to the single-redshift model. We also trained the Emulator without the adversarial loss at both single and multiple redshifts; see Appendix A for details. As shown in Figure 15, we would like to mention the dashed blue lines in the first row, represents the results of the Emulator trained on a single redshift ($z=0$). This single-redshift version shows a significant improvement compared to the multiple-redshift model, reducing the error at the HR Nyquist frequency to around 12%, and the error at $k=10\;h\;{\rm Mpc}^{-1}$ to only 5%. These results represent the best performance the Emulator w/o DG can achieve. However, we have not yet been able to reproduce this level of accuracy when training the Emulator w/o DG across multiple redshifts. While our emulator in its current state is aiming for use cases currently carried out by approximate methods, it is interesting to note that a 5% error in Fourier space at $k=10\;h\;{\rm Mpc}^{-1}$ is comparable to many full N-body simulation codes. (see e.g., Figure 13 of Heitmann et al. 2008).
@@
We can ask what limits the Emulator’s accuracy, and how this can be improved: the movements of particles within a virialized halo are rapid and can vary significantly over time. This high frequency behavior diminishes the Emulator’s ability to accurately match the positions and velocities of particles at the particle level. This difficulty is further compounded by the fact that HRICs do not contain information from non-linear regimes. Additionally, due to memory constraints, the full simulation must be split into cubical chunks, with the GPU handling a maximum of $128^{3}$ particles per chunk. A single large halo may span multiple chunks, meaning the emulator lacks access to all the relevant particle information. This limitation could potentially be mitigated by using a GPU with more memory or use a more memory efficient model. Despite these challenges, the Emulator demonstrates strong performance, effectively adjusting the small-scale details generated by the SR model starting from $k=1\;h\;{\rm Mpc}^{-1}$, achieving excellent accuracy up to the Nyquist frequency of the initial grid. It also maintains decent accuracy up to $k=10\;h\;{\rm Mpc}^{-1}$.
@@
Figure 5: Stochasticity, derived from the correlation coefficient of individual Fourier modes as a function of wavenumber. From left to right, we present results for the cross-correlation of HR (the true N-body simulation) with SR and the Emulator output. The color of each curve corresponds to the redshift, as indicated by the color bar. The shaded area shows the $1\sigma$ standard deviation calculated from all test sets. The vertical dotted gray line is the Nyquist frequency of the LR simulation and the dashed gray line indicates the Nyquist frequency of HR simulation.
@@ 3.3 Halo catalog analysis
Dark matter, under gravitational clustering, forms high-density, virialized structures and local maxima within them, known as halos and subhalos. Understanding these structures and substructures that potentially host galaxies is crucial for connecting theoretical models to observations. These halos and subhalos, identified using halo finder algorithms, play a key role in constructing mock observational catalogs from N-body simulations. We use the SUBFIND code (Springel et al., 2001) to determine the positions and masses of halos and subhalos, as well as to build merger trees. SUBFIND identifies substructures by detecting locally dense and gravitationally bound groups of particles. It starts with the Friends-of-Friends algorithm, followed by the estimation of each particle’s local density via adaptive kernel estimation.
We apply the SUBFIND algorithm to test sets from SR output, Emulator output and HR output of the $N$-body simulation, comparing statistical properties of the halo and subhalo populations. Throughout this analysis, we focus on halos and subhalos containing at least 100 tracer particles in our test sets.
3.3.1 Halo abundance
The most straightforward comparison involves counting the number of halos as a function of mass. The halo mass The halo mass function is defined as:
@@ 3.3.1 Halo abundance
Figure 6 and Figure 7 show the mass function for halos and subhalos measured from the SR, Emulator, and HR outputs at five redshifts from $z=1$ to $z=0$ from the test sets. The colors of the curves represent the redshift of each simulation, with shaded regions indicating the $1\sigma$ standard deviation. In Figure 6, the top panel shows the halo mass function, while the bottom panel presents its fractional error. Similarly, Figure 7 illustrates the subhalo mass function in the same format.
@@
Figure 6: The halo mass function at redshifts $z=1,0.8,0.5,0.2,0$ measured from the SR, Emulator and HR fields. The color of each curve denotes the redshift of simulation in test set. The shaded areas indicate $1\sigma$ standard deviation from all test sets. The upper panel shows halo mass functions and lower panel shows the correspondingfractional errors.
@@
Figure 7: The subhalo mass function at redshifts $z=1,0.8,0.5,0.2,0$ calculated from the SR, Emulator and HR fields. The color of each curve denotes the redshift of simulation in test set. The shaded areas indicate $1\sigma$ standard deviation from all test sets. The upper panel shows halo mass functions and lower panel shows the corresponding fractional errors.
As shown in Figure 6, we observe that the Emulator is biased to lower number densities at lower masses, particularly at $z=0$. This bias arises because HRICs lack information about small mass halos, hence the Emulator can not accurately capturing the correct abundance of low-mass halos. While the combination of adversarial loss helps mitigate this issue, it does not completely resolve it. As shown in Figure 16 ,the absence of guidance from the discriminator exacerbates this bias, resulting in even larger discrepancies.
@@
Regarding substructures, Figure 7 shows that our Emulator model statistically improves over the SR model in the mass range covering $[10^{11},10^{13}]\,M_{\odot}$. The SR model exhibits an approximately 90% deficit in the subhalo mass function at the lower mass end, whereas our Emulator significantly reduces this error by at least 50%.
Figure 8 shows the mean occupation number of subhalos as a function of host halo mass. The SR field exhibits a significant offset from the HR value, predicting $50\%-90\%$ fewer subhalos per halo compared to the HR simulation. The Emulator improves on this from the lower mass end up to $10^{13}\,M_{\odot}$, reducing the error to between 16% and 75%. As shown in Figure 9, for a host halo with a mass of approximately $10^{13}\,M_{\odot}$, the SR output fails to accurately predict the correct number of subhalos. However, the Emulator successfully predicts a subhalo count consistent with the HR simulation and captures their relative positions. Both the SR and Emulator’s outputs struggle to generate sufficient detail, particularly in highly dense regions. As a result, some smaller subhalos are connected by the halo finding algorithm, forming larger subhalos, which leads to an overabundance of massive subhalos and a deficit at the lower-mass end. The Emulator mitigates this overabundance by further refining the SR output. However, one of the very massive subhalo with mass around $10^{14}\,M_{\odot}$ is misclassified as a satellite halo. This misclassification arises because the Emulator’s output remains slightly fuzzier than the target HR simulation, as illustrated in Figure 3 and Figure 2. To address this issue, we manually correct the classification of this halo in our halo catalog.
@@
Both the SR model and the emulator predict fewer subhalos within the most massive halos as shown in Figure 8. We believe that this is partly due to the limited Eulerian field resolution within our discriminator. With the current setup, the input to our discriminator consists of 1024 mesh cells per side of a $100h^{-1}{\rm Mpc}$ box, resulting in a resolution of approximately $0.1h^{-1}{\rm Mpc}$,per Eulerian field "pixel." This resolution limits the discriminator’s ability to accurately detect substructures within massive halos. To quantify the number of subhalos that are not fully resolved, we define the distance between two subhalos as the distance between their centers, subtracting the half-mass radius of both subhalos. We identify subhalos in the HR simulation where the distance to the nearest neighboring subhalo is below the resolution limit (less than the size of one pixel in the Eulerian field). The mean number of such unresolved subhalos per host halo is plotted in the lower Emulator panel of Figure 8 for direct comparison. This average unresolved halo-occupation number reveals that fewer subhalos are fully resolved by the Eulerian field as the host halo mass increases. Increasing the resolution by a factor of $f$ would expand the Eulerian field by a factor of $f^{3}$, demanding more GPU memory. Despite these challenges, our emulator remains quite successful at predicting substructures. Across the halo mass range from $10^{11}-10^{15}\,M_{\odot}$, the mean number of subhalos per halos varies by three orders of magnitude, with the emulator providing relatively accurate results.
@@
We also discover that some halos are divided between simulation chunks that are processed separately. This occurs during both the training and testing phases and causes the model to lack complete halo information. That this occurs is due to GPU memory constraints which requires the simulation to be broken up into relatively small chunks. To quantify the effect of this, we select halos where 99% of their particles are contained within the same chunk. The blue dashed line in the lower panel of Figure 8 the Emulator illustrates the result, showing 10% improvement in the error margin at approximately $10^{11}\,M_{\odot}$. This highlights the significance of this chunking issue in the SR+Emulator procedure. The problem could be mitigated by using a more memory-efficient model or by splitting the model sequentially across multiple GPUs for training and testing.
@@
Figure 8: The mean occupation number of subhalos vs host halo mass at redshift $z=1,0.8,0.5,0.2,0$ measured from the SR, Emulator and HR fields. The $y$ axis is the averaged number of subhalos in each host halo mass bin. The color of each curve corresponds to its redshift according to the color bar at the top. The shaded area shows the $1\sigma$ standard deviation from all test sets. The dashed blue line represents the mean occupation number at $z=0$, calculated by selecting only those halos where 99 percent of their particles lie within a single particle chunk (see text). The solid black line in the lower panel of the Emulator represents the average number of subhalos below the Eulerian resolution of the model, shifted down by 1 for clarity (see text for a detailed description).
@@
Figure 9: Visual comparison of subhalos from SR, Emulator and HR output, centered on a massive halo with halo mass $1.1*10^{13}\,M_{\odot}$. We mark host halos with blue circles and subhalos with red circles. The radius of the circle is proportional to the radius. All halos shown here have mass $>100M_{DM}$.